OpenStudy (anonymous):

Write the given number in the form a + ib 2i^3 - 3i^2 + 5i

OpenStudy (anonymous):

Remember that \(i \cdot i = -1\) What will be \(2 \cdot i^3\) in that case?

OpenStudy (anonymous):

ahhhh 2* i and i^2 will just be -1

OpenStudy (anonymous):

-2*i **** thank you!

OpenStudy (anonymous):

right, so \(2i^3 = -2i\)

OpenStudy (anonymous):

You're welcome =) Now just do for \(-3i^2\) as well and sum it all up